OpenStudy (anonymous):

Solve the following system by substitution. 7x + 2y = 16 –21x – 6y = 24 Please show me step by step how to do this. I used to be pretty good at working with Systems of Equations but I seem to have forgotten how to work this out because I am SO confused.

OpenStudy (texaschic101):

you want to try to get a variable by itself so you can sub it back into the other equation. If you can reduce an equation, do so... for instance , -21x - 6y = 24....divide everything by 3 and you get : -7x - 2y = 8 now we have : -7x - 2y = 8 -- add 2y to both sides -7x = 2y + 8 -- divide by -7 x = -2/7y - 8/7 now sub -2/7y - 8/7 in for x in the other equation 7x + 2y = 16 7(-2/7y - 8/7) + 2y = 16 -- distribute through the parenthesis -2y - 8 + 2y = 16 -- add 8 to both sides -2y + 2y = 16 + 8 0 = 24 (incorrect) there is no solution to this.

OpenStudy (texaschic101):

it is not necessary to reduce the equations...it just makes it easier because you are dealing with smaller numbers do you understand what I did ? Do you have any questions ?

OpenStudy (anonymous):

So when I sub for the variable in the other equation, I'm using the whole new equation for x/y ?

OpenStudy (texaschic101):

um...I do not understand you question. Lets say you have these equations.. x + 3y = 5 ----> x = -3y + 5 2x + 4y = 7 now you would sub -3y + 5 in for x in the other equation...like this 2(-3y + 5) + 4y = 7 does that make sense to you ?

OpenStudy (texaschic101):

let me do it without reducing.... 7x + 2y = 16 2y = -7x + 16 y = -7/2x + 8 now sub -7/2x + 8 in for y in the other equation -21x - 6y = 24 -21x - 6(-7/2x + 8) = 24 -21x + 21x - 48 = 24 -21x + 21x = 24 + 48 -21x + 21x = 72 0 = 72 (incorrect) NO SOLUTION

OpenStudy (anonymous):

You answered my questions. Thank you very much; I think I understand. ♥

OpenStudy (texaschic101):

good to hear :)